# from perception to punchline: empowering vlm with the art of in-the-wild meme

### v1

@@ Abstract
Generating humorous memes is a challenging multimodal task that moves beyond direct image-to-caption supervision. It requires a nuanced reasoning over visual content, contextual cues, and subjective humor. To bridge this gap between visual perception and humorous punchline creation, we propose HUMOR, a novel framework that guides VLMs through hierarchical reasoning and aligns them with group-wise human preferences. First, HUMOR employs a hierarchical, multi-path Chain-of-Thought (CoT): the model begins by identifying a template-level intent, then explores diverse reasoning paths under different contexts, and finally anchors onto a high-quality, context-specific path. This CoT supervision, which traces back from ground-truth captions, enhances reasoning diversity. We further analyze that this multi-path exploration with anchoring maintains a high expected humor quality, under the practical condition that high-quality paths retain significant probability mass. Second, to capture subjective humor, we train a pairwise reward model that operates within groups of memes sharing the same template. Following established theory, this approach ensures a consistent and robust proxy for human preference, even with subjective and noisy labels. The reward model then enables a group-wise reinforcement learning optimization, guaranteeing providing a theoretical guarantee for monotonic improvement within the trust region. Extensive experiments show that HUMOR empowers various VLMs with superior reasoning diversity, more reliable preference alignment, and higher overall meme quality. Beyond memes, our work presents a general training paradigm for open-ended, human-aligned multimodal generation, where success is guided by comparative judgment within coherent output groups.
@@ 1 Introduction
Creativity in multimodal understanding and generation is progressively shifting from literal description and visual perception to subjective and context-dependent outcomes, such as humor (hessel2022androids; hwang2023memecap), aesthetics (rombach2022high), and social alignment (bai2022constitutional). In these domains, quality is not defined by a single ground-truth but is instead guided by diverse—and often noisy—distributions of human preference (yadav2025beyond; burn2018multimodality; song2025large). While recent vision–language models (VLMs) have scaled to achieve strong results on captioning and visual question answering (kuang2025natural; ghandi2023deep), which admit relatively objective targets (yan2023achieving), the question of how to align models with such open-ended values (bhatia2024local; feng2024modular) remains an open challenge.
As a salient instance and testbed of this open-ended challenge, meme punchline generation illuminates these limitations. Current approaches often treat it as a straightforward image-to-caption task, optimized with fixed supervised losses on collected data (peirson2018dank; vyas2023pro; hwang2023memecap). This formulation collapses the inherently diverse reasoning required for humor into the language model, suppresses intermediate interpretability, and tends to yield outputs that are fluent yet shallow or unfunny (yadav2025beyond). Direct generation from a template forfeits control over the interpretive process and makes targeted steering difficult. Here a template denotes the shared visual basis or theme image for a group of memes. While recent evidence shows chain-of-thought (CoT) reasoning improves performance in VLMs (zhang2023multimodal; hu2024visual; lad2025), we argue that meme generation necessitates a hierarchical, multi-path reasoning process. Observations from real-world meme data suggest that multiple reasoning paths can lead to distinct metaphorical bindings or punchlines (xu2022met). Concretely, we decompose the process into two stages: a template-level stage that infers a canonical intent, followed by a context-level stage that grounds this intent in specific visual details. By first exploring multiple reasoning paths and then explicitly anchoring one with supervision, this design ensure diversity and maintain enough humor quality. Exploring multiple reasoning paths—and subsequently anchoring one with supervision—ensures diversity. As our theoretical analysis (Section 4.3) shows, this approach preserves a conditional lower bound on expected humor, provided high-quality paths retain sufficient probability mass and remaining paths are not significantly worse.
@@
Furthermore, optimizing generation toward human-preferred humor is essential to meet these theoretical conditions. Prior work on computational humor often relies on text-only cues or global funniness scores  (baluja2024text; kalloniatis2024computational; zhu2025commonality), implicitly assuming humor is directly comparable across different meme templates. In practice, however, human judgments are more reliable and consistent within a coherent template group than across groups with differing conventions. Ignoring this group structure introduces noise, harms generalization, and encourages models to exploit reward superficial overlaps rather than genuine humorous fit(see empirical analyses in Sec. 5.2 and Sec. 5.4). Although humor cannot be reduced to a deterministic rule, we can design a pairwise reward model that enforces rank consistency within each template group. This model provides a stable, human-aligned preference signal, and further enables group-wise RL to ensure that expected humor cannot degrade beyond a bounded amount.
@@
    Comprehensive experiments across multiple models and data demonstrate that HUMOR improves reasoning diversity, preference alignment, and overall meme quality.
@@ 2.1 Evolution of Vision-Language Models for Multi-modal Process
The pursuit of unified vision-language modeling has progressed through three distinct phases of architectural innovation. Early foundational work established bidirectional frameworks for cross-modal understanding: ERNIE-ViLG (Zhang2021ERNIEViLGUG) and the Unifying Multi-modal Transformer (10.1145/3474085.3481540) pioneered transformer-based architectures that jointly optimized text-to-image and image-to-text generation through multi-modal tokenization and autoregressive objectives. Concurrently, Zero-Shot Text-to-Image Generation (Ramesh2021ZeroShotTG) demonstrated the scalability potential of such approaches through their zero-shot text-to-image generation framework, establishing critical baselines for large-scale multi-modal pretraining.
Contemporary breakthroughs have redefined architectural paradigms through multimodal unification. Models like Show-o (xie2024show) and MonoFormer (zhao2024monoformer) successfully fused autoregressive and diffusion mechanisms within singular architectures via shared attention layers. Beyond architectural fusion, recent research highlights the critical role of reasoning strategies. Chain-of-Thought (CoT) prompting has been empirically shown to enhance the complex reasoning capabilities of VLMs by eliciting intermediate rationales (zhang2023multimodal; hu2024visual). Building upon these advancements, our work leverages multi-modal comprehension capabilities to address the unique challenges of meme generation—particularly its requirement for hierarchical reasoning and understanding subjective humor.
@@ 2.2 Meme Analysis, Generation, and Alignment
Computational humor draws from established linguistic and anthropological theories (apte1985humor; binsted2006computational) to formally model incongruity and semantic shifts. Internet memes have emerged as a vital component of digital culture, prompting substantial scholarly attention to their multi-modal communications. Extensive research has focused on analyzing topics (du2020understanding), semantics (xu2022met), and emotions (sharma2020semeval) conveyed in memes. The evolution of meme generation techniques has progressed through distinct technological phases. Initial systems employed rule-based architectures, exemplified by  oliveira2016one’s template-driven approach using standardized structures like ”One does not simply X”, and  wang2015can’s dual-channel model integrating textual and visual features. The advent of deep learning catalyzed more sophisticated paradigms. Peirson and Tolunay pioneered this transition with Dank Learning (peirson2018dank), combining Inception V3 image encoders with attention-enhanced LSTM decoders. Subsequent innovations introduced transformer architectures: Sadasivam et al.’s MemeBot (sadasivam2020memebot) and Vyalla et al.’s Memeify (vyalla2020memeify) demonstrated enhanced text-image alignment through multi-modal fusion.
Recent breakthroughs leverage large language models (LLMs) and VLMs to achieve unprecedented scale. Memecraft (wang2024memecraft) enables targeted meme creation for social advocacy. Addressing multi-image complexity, Chen et al. proposed XMeCap (chen2024xmecap), introducing a two-stage framework with supervised fine-tuning guided by novel similarity metrics. Concurrently, benchmark datasets have emerged to evaluate capabilities. MemeCap (hwang2023memecap) provides metaphor annotations, while the New Yorker benchmarks series (hessel2022androids; hessel2023winning) assess humor comprehension through caption matching and explanation tasks. The MCC dataset (sharma2023memex) further incorporates external knowledge for abstraction analysis.
While capability has scaled, aligning models with subjective human preferences remains a critical frontier. Unlike objective tasks with ground-truth, humor and creativity require modeling diverse and often noisy judgments. Recent works have begun to address this by aligning models with diverse human values (zhou2024beyond) and exploring personalized or pluralistic strategies (feng2024modular). Specifically in the domain of humor, song2025large highlight the challenges of modeling subjective humor preferences using LLMs. Our work advances this direction by proposing a group-wise preference formulation, mitigating the noise inherent in cross-context humor comparison.
@@ 3 Problem Formulation
For a given group $G$, we collect human annotations indicating which of two memes is considered funnier. Formally, for $m_{i},m_{j}\in G$, define $y_{ij}^{G}\;=\;\mathbb{I}[\,m_{i}\succ m_{j}\,]\in\{0,1\}$ where $m_{i}\succ m_{j}$ denotes a local preference that $m_{i}$ is judged to be funnier than $m_{j}$. The dataset consists of triples $(G,(m_{i},m_{j}),y_{ij}^{G})$ sampled from a pairing distribution over $G$. We allow for incompleteness (not all pairs are labeled) and noise (due to inter-annotator disagreement). We adopt two mild yet standard assumptions from preference learning christiano2023deepreinforcementlearninghuman: (i) local comparability: preferences are elicited and interpreted only within a fixed group $G$; (ii) weak transitivity: in expectation, if $m_{i}\succ m_{j}$ and $m_{j}\succ m_{\ell}$, then $m_{i}\succ m_{\ell}$ is more likely than its reversal, without requiring a strict total order.
@@
where $\Lambda:\mathbb{R}\!\to\!(0,1)$ is a strictly increasing link function (e.g., logistic or probit)  (sun2025rethinking). Eq. 1 captures the intuition that the probability of preferring $m_{i}$ to $m_{j}$ depends only on their latent humor gap within the same group: when $h_{G}(m_{i})\!\approx\!h_{G}(m_{j})$, the choice is nearly ambiguous (probability $\approx\!1/2$); as the gap increases, the probability moves smoothly toward $1$ (if $h_{G}(m_{i})\!>\!h_{G}(m_{j})$) or 0 (otherwise), capturing that larger humor gaps lead to more consistent comparisons.
@@ 4.1 Hierarchical Chain-of-Thought Supervision
Meme creation mirrors a hierarchical cognitive routine: humans first parse what a visual template affords, and then realize a chosen intent with text that fits the surrounding context flamson2008encryption. We therefore model meme generation as a two-stage reasoning process, separating (i) intent inference from the image and (ii) context-sensitive textual realization of that intent. In practice, however, training trajectories are often single-path because they are derived from a single gold caption: back-deriving a rationale from one answer yields only one route.
@@ 4.2 Reward Modeling from Pairwise Preferences
The ideal learning objective would be to recover the latent humor function $h_{G}(m)$ for each meme $m$. Since humor is inherently subjective and lacks a global scale, this is infeasible in practice. We therefore adopt an order-consistent view of reward modeling (following established theory (sun2025rethinking)) and instantiate it in our group-wise meme setting. Under this formulation, the reward serves as a within-group surrogate of $h_{G}$, trained only from relative judgments, avoiding ill-posed cross-group calibration. Intuitively, hierarchical CoT ensures that high-quality paths retain a meaningful probability mass (the $\alpha$ condition via Stage 2 anchoring), while the reward model supplies the preference signal necessary to shrink the average gap among plausible paths (addressing the $\delta$ condition). This transforms open-ended exploration into a tractable selection problem.
@@
These propositions, while grounded in the order-consistent analysis of sun2025rethinking, are specifically instantiated under our group-wise comparability. They serve as the theoretical drivers to reduce the humor gap $\delta$ after CoT has secured $\alpha$. Since pairwise data can be sparse, we aggregate $\widehat{p}^{\,G}_{ij}$ into a coherent within-group ranking via Expected Borda Count (EBC) (see Appendix D for more explanations and implementations). For a candidate set $\mathcal{S}_{G}$, each meme’s EBC score equals its expected number of wins against others under the model in Eq. 3. This provides a stable training target, and inherits expected order consistency when the pairwise model is consistent (Appendix B). To complement this outcome-oriented supervision, we further incorporate auxiliary rewards to regulate the intermediate reasoning path (Appendix F), comprising a structural format reward (Appendix F.1) and a monotonicity-validated content reward (Appendix F.2). Detailed constructs and examples of pairwise data are provided in Appendix E.
@@ 4.3 Group-wise Policy Optimization
Following the CoT supervision stage and reward model training, we further fine-tune the meme generator to increase the probability of higher-ranked captions. Concretely, we leverage the trained reward model and adopt a Group-wise Relative Policy Optimization (GRPO) objective shao2024deepseekmathpushinglimitsmathematical. For a candidate set of memes $\mathcal{S}_{G}$ with ranking $q_{G}$ from EBC, the reinforcement fine-tuning loss is:
@@
where $\pi_{\text{ref}}$ denotes the policy obtained after CoT training. The first term aligns $\pi_{\theta}$ with the group-local preference distribution $q_{G}$ (rank-consistent with $h_{G}$). Since prior perference optimization analyses  christiano2023deepreinforcementlearninghuman; neu2012apprenticeshiplearningusinginverse; haarnoja2018softactorcriticoffpolicymaximum often propose optimistic lower bounds (second term), we also adopt a corrected, KL-controlled guarantee that holds under our setting and noise model. Specifically, the original upper bound on the humor-score deviation (induced by preference noise) can be refined under GRPO into a bound that scales with the KL between the trained policy and the reference policy (proof in Appendix C).

### v2

@@ Abstract
Generating humorous memes is a challenging multimodal task that moves beyond direct image-to-caption supervision. It requires a nuanced reasoning over visual content, contextual cues, and subjective humor. To bridge this gap between visual perception and humorous punchline creation, we propose HUMOR, a novel framework that guides VLMs through hierarchical reasoning and aligns them with group-wise human preferences. First, HUMOR employs a hierarchical, multi-path Chain-of-Thought (CoT): the model begins by identifying a template-level intent, then explores diverse reasoning paths under different contexts, and finally anchors onto a high-quality, context-specific path. This CoT supervision, which traces back from ground-truth captions, enhances reasoning diversity. We further analyze that this multi-path exploration with anchoring maintains a high expected humor quality, under the practical condition that high-quality paths retain significant probability mass. Second, to capture subjective humor, we train a pairwise reward model that operates within groups of memes sharing the same template. Following established theory, this approach ensures a consistent and robust proxy for human preference, even with subjective and noisy labels. The reward model then enables a group-wise reinforcement learning optimization, guaranteeing providing a theoretical guarantee for monotonic improvement within the trust region. Extensive experiments show that HUMOR empowers various VLMs with superior reasoning diversity, more reliable preference alignment, and higher overall meme quality. Beyond memes, our work presents a general training paradigm for open-ended, human-aligned multimodal generation, where success is guided by comparative judgment within coherent output groups. Our models are publicly available at https://huggingface.co/OpenDILabCommunity/HUMOR-COT-Qwen2.5-VL and https://huggingface.co/OpenDILabCommunity/HUMOR-RM-Keye-VL.
@@ 1 Introduction
Creativity in multimodal understanding and generation is progressively shifting from literal description and visual perception to subjective and context-dependent outcomes, such as humor (Hessel et al., 2023; Hwang and Shwartz, 2023), aesthetics (Rombach et al., 2022), and social alignment (Bai et al., 2022). In these domains, quality is not defined by a single ground-truth but is instead guided by diverse—and often noisy—distributions of human preference (Yadav et al., 2025; Burn and Kress, 2018; Song et al., 2025). While recent vision–language models (VLMs) have scaled to achieve strong results on captioning and visual question answering (Kuang et al., 2025; Ghandi et al., 2023), which admit relatively objective targets (Yan et al., 2023), the question of how to align models with such open-ended values (Bhatia et al., 2024; Feng et al., 2024) remains an open challenge.
As a salient instance and testbed of this open-ended challenge, meme punchline generation illuminates these limitations. Current approaches often treat it as a straightforward image-to-caption task, optimized with fixed supervised losses on collected data (Peirson V and Tolunay, 2018; Vyas et al., 2023; Hwang and Shwartz, 2023). This formulation collapses the inherently diverse reasoning required for humor into the language model, suppresses intermediate interpretability, and tends to yield outputs that are fluent yet shallow or unfunny (Yadav et al., 2025). Direct generation from a template forfeits control over the interpretive process and makes targeted steering difficult. Here a template denotes the shared visual basis or theme image for a group of memes. While recent evidence shows chain-of-thought (CoT) reasoning improves performance in VLMs (Zhang et al., 2023; Hu et al., 2024; Zhang and Niu, 2025), we argue that meme generation necessitates a hierarchical, multi-path reasoning process. Observations from real-world meme data suggest that multiple reasoning paths can lead to distinct metaphorical bindings or punchlines (Xu et al., 2022). Concretely, we decompose the process into two stages: a template-level stage that infers a canonical intent, followed by a context-level stage that grounds this intent in specific visual details. By first exploring multiple reasoning paths and then explicitly anchoring one with supervision, this design ensure diversity and maintain enough humor quality. Exploring multiple reasoning paths—and subsequently anchoring one with supervision—ensures diversity. As our theoretical analysis (Section 4.3) shows, this approach preserves a conditional lower bound on expected humor, provided high-quality paths retain sufficient probability mass and remaining paths are not significantly worse.
@@
Furthermore, optimizing generation toward human-preferred humor is essential to meet these theoretical conditions. Prior work on computational humor often relies on text-only cues or global funniness scores  (Baluja, 2024; Kalloniatis and Adamidis, 2024; Zhu et al., 2025a), implicitly assuming humor is directly comparable across different meme templates. In practice, however, human judgments are more reliable and consistent within a coherent template group than across groups with differing conventions. Ignoring this group structure introduces noise, harms generalization, and encourages models to exploit reward superficial overlaps rather than genuine humorous fit(see empirical analyses in Sec. 5.2 and Sec. 5.4). Although humor cannot be reduced to a deterministic rule, we can design a pairwise reward model that enforces rank consistency within each template group. This model provides a stable, human-aligned preference signal, and further enables group-wise RL to ensure that expected humor cannot degrade beyond a bounded amount.
@@
    Comprehensive experiments across multiple models and data demonstrate that HUMOR improves reasoning diversity, preference alignment, and overall meme quality. To facilitate future research, we have released our HUMOR-COT and HUMOR-RM models.
@@ 2.1 Evolution of Vision-Language Models for Multi-modal Process
The pursuit of unified vision-language modeling has progressed through three distinct phases of architectural innovation. Early foundational work established bidirectional frameworks for cross-modal understanding: ERNIE-ViLG (Zhang et al., 2021) and the Unifying Multi-modal Transformer (Huang et al., 2021) pioneered transformer-based architectures that jointly optimized text-to-image and image-to-text generation through multi-modal tokenization and autoregressive objectives. Concurrently, Zero-Shot Text-to-Image Generation (Ramesh et al., 2021) demonstrated the scalability potential of such approaches through their zero-shot text-to-image generation framework, establishing critical baselines for large-scale multi-modal pretraining.
Contemporary breakthroughs have redefined architectural paradigms through multimodal unification. Models like Show-o (Xie et al., 2024) and MonoFormer (Zhao et al., 2024) successfully fused autoregressive and diffusion mechanisms within singular architectures via shared attention layers. Beyond architectural fusion, recent research highlights the critical role of reasoning strategies. Chain-of-Thought (CoT) prompting has been empirically shown to enhance the complex reasoning capabilities of VLMs by eliciting intermediate rationales (Zhang et al., 2023; Hu et al., 2024). Building upon these advancements, our work leverages multi-modal comprehension capabilities to address the unique challenges of meme generation—particularly its requirement for hierarchical reasoning and understanding subjective humor.
@@ 2.2 Meme Analysis, Generation, and Alignment
Computational humor draws from established linguistic and anthropological theories (Apte, 1985; Binsted et al., 2006) to formally model incongruity and semantic shifts. Internet memes have emerged as a vital component of digital culture, prompting substantial scholarly attention to their multi-modal communications. Extensive research has focused on analyzing topics (Du et al., 2020), semantics (Xu et al., 2022), and emotions (Sharma et al., 2020) conveyed in memes. The evolution of meme generation techniques has progressed through distinct technological phases. Initial systems employed rule-based architectures, exemplified by  Oliveira et al. (2016)’s template-driven approach using standardized structures like ”One does not simply X”, and  Wang and Wen (2015)’s dual-channel model integrating textual and visual features. The advent of deep learning catalyzed more sophisticated paradigms. Peirson and Tolunay pioneered this transition with Dank Learning (Peirson V and Tolunay, 2018), combining Inception V3 image encoders with attention-enhanced LSTM decoders. Subsequent innovations introduced transformer architectures: Sadasivam et al.’s MemeBot (Sadasivam et al., 2020) and Vyalla et al.’s Memeify (Vyalla and Udandarao, 2020) demonstrated enhanced text-image alignment through multi-modal fusion.
Recent breakthroughs leverage large language models (LLMs) and VLMs to achieve unprecedented scale. Memecraft (Wang and Lee, 2024) enables targeted meme creation for social advocacy. Addressing multi-image complexity, Chen et al. proposed XMeCap (Chen et al., 2024b), introducing a two-stage framework with supervised fine-tuning guided by novel similarity metrics. Concurrently, benchmark datasets have emerged to evaluate capabilities. MemeCap (Hwang and Shwartz, 2023) provides metaphor annotations, while the New Yorker benchmarks series (Hessel et al., 2023) assess humor comprehension through caption matching and explanation tasks. The MCC dataset (Sharma et al., 2023) further incorporates external knowledge for abstraction analysis.
While capability has scaled, aligning models with subjective human preferences remains a critical frontier. Unlike objective tasks with ground-truth, humor and creativity require modeling diverse and often noisy judgments. Recent works have begun to address this by aligning models with diverse human values (Zhou et al., 2024) and exploring personalized or pluralistic strategies (Feng et al., 2024). Specifically in the domain of humor, Song et al. (2025) highlight the challenges of modeling subjective humor preferences using LLMs. Our work advances this direction by proposing a group-wise preference formulation, mitigating the noise inherent in cross-context humor comparison.
@@ 3 Problem Formulation
For a given group $G$, we collect human annotations indicating which of two memes is considered funnier. Formally, for $m_{i},m_{j}\in G$, define $y_{ij}^{G}\;=\;\mathbb{I}[\,m_{i}\succ m_{j}\,]\in\{0,1\}$ where $m_{i}\succ m_{j}$ denotes a local preference that $m_{i}$ is judged to be funnier than $m_{j}$. The dataset consists of triples $(G,(m_{i},m_{j}),y_{ij}^{G})$ sampled from a pairing distribution over $G$. We allow for incompleteness (not all pairs are labeled) and noise (due to inter-annotator disagreement). We adopt two mild yet standard assumptions from preference learning Christiano et al. (2023): (i) local comparability: preferences are elicited and interpreted only within a fixed group $G$; (ii) weak transitivity: in expectation, if $m_{i}\succ m_{j}$ and $m_{j}\succ m_{\ell}$, then $m_{i}\succ m_{\ell}$ is more likely than its reversal, without requiring a strict total order.
@@
where $\Lambda:\mathbb{R}\!\to\!(0,1)$ is a strictly increasing link function (e.g., logistic or probit)  (Sun et al., 2025). Eq. 1 captures the intuition that the probability of preferring $m_{i}$ to $m_{j}$ depends only on their latent humor gap within the same group: when $h_{G}(m_{i})\!\approx\!h_{G}(m_{j})$, the choice is nearly ambiguous (probability $\approx\!1/2$); as the gap increases, the probability moves smoothly toward $1$ (if $h_{G}(m_{i})\!>\!h_{G}(m_{j})$) or 0 (otherwise), capturing that larger humor gaps lead to more consistent comparisons.
@@ 4.1 Hierarchical Chain-of-Thought Supervision
Meme creation mirrors a hierarchical cognitive routine: humans first parse what a visual template affords, and then realize a chosen intent with text that fits the surrounding context Flamson and Barrett (2008). We therefore model meme generation as a two-stage reasoning process, separating (i) intent inference from the image and (ii) context-sensitive textual realization of that intent. In practice, however, training trajectories are often single-path because they are derived from a single gold caption: back-deriving a rationale from one answer yields only one route.
@@ 4.2 Reward Modeling from Pairwise Preferences
The ideal learning objective would be to recover the latent humor function $h_{G}(m)$ for each meme $m$. Since humor is inherently subjective and lacks a global scale, this is infeasible in practice. We therefore adopt an order-consistent view of reward modeling (following established theory (Sun et al., 2025)) and instantiate it in our group-wise meme setting. Under this formulation, the reward serves as a within-group surrogate of $h_{G}$, trained only from relative judgments, avoiding ill-posed cross-group calibration. Intuitively, hierarchical CoT ensures that high-quality paths retain a meaningful probability mass (the $\alpha$ condition via Stage 2 anchoring), while the reward model supplies the preference signal necessary to shrink the average gap among plausible paths (addressing the $\delta$ condition). This transforms open-ended exploration into a tractable selection problem.
@@
These propositions, while grounded in the order-consistent analysis of Sun et al. (2025), are specifically instantiated under our group-wise comparability. They serve as the theoretical drivers to reduce the humor gap $\delta$ after CoT has secured $\alpha$. Since pairwise data can be sparse, we aggregate $\widehat{p}^{\,G}_{ij}$ into a coherent within-group ranking via Expected Borda Count (EBC) (see Appendix D for more explanations and implementations). For a candidate set $\mathcal{S}_{G}$, each meme’s EBC score equals its expected number of wins against others under the model in Eq. 3. This provides a stable training target, and inherits expected order consistency when the pairwise model is consistent (Appendix B). To complement this outcome-oriented supervision, we further incorporate auxiliary rewards to regulate the intermediate reasoning path (Appendix F), comprising a structural format reward (Appendix F.1) and a monotonicity-validated content reward (Appendix F.2). Detailed constructs and examples of pairwise data are provided in Appendix E.
@@ 4.3 Group-wise Policy Optimization
Following the CoT supervision stage and reward model training, we further fine-tune the meme generator to increase the probability of higher-ranked captions. Concretely, we leverage the trained reward model and adopt a Group-wise Relative Policy Optimization (GRPO) objective Shao et al. (2024). For a candidate set of memes $\mathcal{S}_{G}$ with ranking $q_{G}$ from EBC, the reinforcement fine-tuning loss is:
@@
where $\pi_{\text{ref}}$ denotes the policy obtained after CoT training. The first term aligns $\pi_{\theta}$ with the group-local preference distribution $q_{G}$ (rank-consistent with $h_{G}$). Since prior perference optimization analyses  Christiano et al. (2023); Neu and Szepesvari (2012); Haarnoja et al. (2018) often propose optimistic lower bounds (second term), we also adopt a corrected, KL-controlled guarantee that holds under our setting and noise model. Specifically, the original upper bound on the humor-score deviation (induced by preference noise) can be refined under GRPO into a bound that scales with the KL between the trained policy and the reference policy (proof in Appendix C).